### v1

@@ 5.2.2 Hyperspace Transformation
3:     $S_{next}=SelectNext(M,S_{best})$ /* use EI acquisition function to select new sample points on the surrogate model $M$*/
4:     $R_{next}=ReConstruct(S_{next})$
5:     $Score_{next}=BatchEval(S_{next},D,Q)$
6:     Update $M$ based on $(S_{next},Score_{next})$
[…]
8:        $Score_{best}=\text{max}(Score_{next})$
9:        $S_{best}=argmax(Score_{next})$
10:        $R_{best}=R_{next}$
11:     end if
12:  end for
13:  return  $R_{best},S_{best}$
Algorithm 1 Matrix Optimization
@@ 5.2.3 Hyperspace Movement
Hyperspace movement further optimizes the feature vectors on the result of previous hyperspace transformation by moving similar data points in hyperspace together to get $\hat{D}$, where similar data can be placed in adjacent places or the same cluster. Thus it could benefit index building and multimodal data retrieval.
Figure 10: Hyperspace Movement
To achieve this objective, we propose a new data movement method called Local Parallelized Gravitational Field (LPGF) which improves on an existing data movement method HIBOG (Li et al., 2021). HIBOG posits that in hyperspace, a data point is attracted by its similar data points, causing it to move towards the resultant force direction of these attractions. Specifically, it defines a gravitation force between each data point and its nearest $K$ points and applies this force to this data point to make displacement. Though HIBOG shows effectiveness in gathering similar data, it suffers the following drawbacks: (1) Computationally expensive: Finding the nearest $K$ data points for force calculation requires sorting all data points, which requires huge computation, especially for large datasets. Additionally, the parameter $K$ needs to be set and adjusted, further increasing the computation time. (2) Anomalies of movement: The calculation of forces between data points may not be suitable for all scenarios. Sometimes, this force can cause anomalies in the movement of tightly clustered data. (3) Weak parallel capability: The spatial range of the nearest $K$ neighbors is not predetermined, leading to the inability to perform grid partitioning and parallel processing of data. To overcome these shortcomings, we introduce the pipeline of LPGF, which consists of three steps to improve the above shortcomings, as shown in Fig 10.
@@
Figure 11: Force Calculation Formula
@@ 5.2.4 Feature Representation Evaluation
To prove that our method could successfully improve the performance of clustering within the dataset, which forms a solid foundation for index construction, we design experiments to demonstrate the improvement of clustering (Evaluation One) and high-dimensional learned index query performance (Evaluation Two) by our proposed feature representation method.
Evaluation One:  Clustering Enhancement Evaluation
To validate the capability of our proposed methods to gather similar multimodal data together, we compare five feature representation methods for enhancing clustering performance, which are T, HIBOG, LPGF, T + HIBOG, and T + LPGF. T represents Hyperspace Transformation. HIBOG and LPGF denote Hyperspace Movement. T + HIBOG as well as T + LPGF represent the combination of Hyperspace Transformation and Hyperspace Movement. The experiments are conducted using the real dataset Flame (Fu and Medico, 2007) and employ three well-known clustering methods, which are K-means, Agglomerative (Sokal and Michener, 1958), and Density Peaks Clustering (DPC) (Rodriguez and Laio, 2014), to represent diverse clustering paradigms. The selection of these methods is not meant to imply exclusivity but rather to demonstrate the generalization ability across various clustering algorithms. To comprehensively evaluate the impact of different representation methods on clustering performance, we use the Silhouette Coefficient, Calinski-Harabasz Index, and Normalized Mutual Information as experimental metrics. These metrics assess clustering effectiveness from different perspectives.
@@
The experimental results of different feature representation methods are presented in Table 5. Comparing the results of LPGF and HIBOG, we observe that LPGF significantly enhances clustering performance across various algorithms, suggesting that optimizations in gravitational field design in LPGF improve intra-class compactness and inter-class separation more effectively than HIBOG. Additionally, T + HIBOG and T + LPGF demonstrate that the hyperspace transformation matrix T not only enhances the discriminative features of multimodal data but also substantially improves clustering effectiveness.
@@
Evaluation Two:  High-dimensional Learned Index Query Performance Enhancement Evaluation
We design two experiments to demonstrate the effectiveness of our proposed feature representation method in enhancing the performance of high-dimensional learned index queries. The first experiment demonstrates that our method can achieve excellent "last-mile" models within the leaf nodes, while the second evaluates their influence on query performance of currently typical high-dimensional learned indexes. The experiments utilize a generated dataset comprising 2M records, with data distributed according to a Gaussian mixture distribution.
@@
In the first experiment, we compare the smoothness of Cumulative Distribution Function (CDF) curves of datasets processed by different feature representation methods. Smooth CDF curves help address the "last mile problem" in learned index, making it easier for simple machine learning and improving the accuracy of index predictions. We compare the CDF curves of four datasets, which are Original Dataset, HIBOG Optimized Dataset, LPGF Optimized Dataset and T + LPGF Optimized Dataset. As shown in Fig 12, by comparing the smoothness of the CDF curves for each dataset, we find that T + LPGF achieves the highest improvement in the smoothness among all optimized methods. Compared to HIBOG, LPGF significantly smoothens the global CDF curve, because the force calculation method in LPGF reduces the probability of anomalous data movement, resulting in tighter data clustering. T + LPGF not only retains the advantages of LPGF but also highlights critical clustering features through matrix T, further enhancing the clustering capabilities. Overall, the T + LPGF method results in tighter multimodal data clustering, improving the smoothness of the CDF curve, and effectively reducing the complexity of the machine learning involved in index construction.
Figure 12: The CDF Curves of Different Feature Representation Methods
Figure 13: The Average Query Times of Different Indexes on Original Dataset and Optimized Dataset
@@
In the second experiment, we demonstrate the effectiveness of our methods in improving the efficiency of typical high-dimensional learned indexes by comparing the average query times. We select five typical high-dimensional learned indexes, which are ML, ZM, LISA, Flood, and LIMS. These learned indexes all share the characteristic that their query performance is significantly influenced by the smoothness of the CDF curve which directly impacts "last-mile" searching. As shown in Fig 13, we compare the average query times of different indexes on the original, the LPGF optimized, and the T + LPGF optimized datasets. The experimental results indicate that both LPGF and T + LPGF methods significantly improve the query performance, with T + LPGF performing the best. ML and LIMS show the most significant performance improvements across all datasets, because the creation of these indexes relies heavily on clustering results, confirming the effectiveness of the T + LPGF method in enhancing data clustering. LISA and Flood, being variants of grid indexes, benefit from the ability of T + LPGF to make data more tightly clustered, thereby improving the smoothness of the CDF curve within the grid. Since the calculation of z-addresses in ZM is closely related to the positions of data in hyperspace, a tighter data distribution can improve its continuity, making the data easier for machine learning. The T + LPGF method improves the overall distribution and continuity of data, providing smoother CDF curves that are easier for various indexes to learn, effectively enhancing the query performance of high-dimensional learned indexes.
@@ 6 High-dimensional Learned Index
To fully leverage the advantages of the "Multimodal Data Feature Representation" introduced in Section 5 and further enhance multimodal data retrieval efficiency, we propose a high-dimensional learned index. This index hierarchically organizes clusters and their sub-clusters to form a cluster tree, where the index of the leaf nodes is established through a "last-mile" training approach. Furthermore, the index’s inner structure can be optimized through our query-aware mechanism. An overview of our index is illustrated in Fig 14, detailing both the construction and optimization processes. During the construction phase, divisive hierarchical clustering is employed to form clusters at various levels, corresponding to different levels within the tree structure, with the leaf nodes representing the most granular, indivisible sub-clusters, thereby substantially enhancing the precision of the "last-mile" training for the index. Concurrently, the optimization phase employs a query-aware mechanism to refine the internal structure of the index, improving its adaptability to diverse query demands. The detailed processes of index construction and optimization are elaborated in Section 6.1 and Section 6.2, respectively.
Figure 14: Overview of Index Construction

### v2

@@ 5.2.2 Hyperspace Transformation
3:     Fit local gaussian process models $LGP=\{lgp_{1},...lgp_{n}\}$
4:     $(R_{next}^{t},S_{next}^{t})_{t\in[1,n]}=SelectNext(LGP,T)$ /*Select new candidates in each trust region*/
5:     $((Y_{time}^{t},Y_{CBR}^{t},Y_{acc}^{t})_{t\in[1,n]})=\text{BatchEval}(f,D,Q;(R_{t}^{next},S_{t}^{next})_{t\in[1,n]}$
6:     $X_{0}.append((R_{next}^{t},S_{next}^{t})_{t\in[1,n]})$,$Y_{0}.append((Y_{time}^{t},Y_{CBR}^{t},Y_{acc}^{t})_{t\in[1,n]})$
7:     for j in [0, n] do
8:        $T_{j},L_{j}=Updata(T_{j},(R_{next}^{j},S_{next^{\prime}}^{j}Y_{time^{\prime}}^{j}Y_{CBR^{\prime}}^{j}Y_{acc}^{j}))$
9:        if $L_{j}<L_{\min}$ then
10:           Terminate $T_{j}$
11:           $T_{j},X_{0},Y_{0}=Reinitialize(L_{init},X_{0},Y_{0})$ /*Add new region centroid and objective function values into $X_{0}$, $Y_{0}$*/
12:        else
13:           update $T_{j}$’s region center
14:        end if
15:     end for
16:  end for
17:  $P=SelectPF(X_{k},Y_{k})$
18:  return  $P$
Algorithm 1 Matrix Optimization Based On MORBO
Hyperspace Transformation Evaluation
Figure 10: The cost time of T construction and data transformation with different number of tasks and dataset sizes. (a)Dataset size = 100M, (b) Task Number = 40.
The hyperspace transformation process in MQRLD is executed offline, utilizing large-scale of historical data. To provide a comprehensive analysis of this process, we evaluate both the complexity and the effectiveness of hyperspace transformation. The complexity is assessed through theoretical and experimental analysis, and the effectiveness is assessed by comparing hyperspace transformation with other similar approaches.
First, we evaluate the complexity of hyperspace transformation. In Step 1, the matrix $D$ is obtained, $D\in\mathbb{R}^{N\times d}$, where $N$ is the number of data points, ranging from 0.30M to 210M in our experiments, and $d$ is the dimensionality of each data points, ranging from 3 to 1026. Storing the matrix $D$ requires space complexity of $O(N\times d)$. Step 2 involves the computation and storage of the covariance matrix $C$, which has a time complexity of $O(N\times d^{2})$ and a space complexity of $O(d^{2})$. In Step 3, the matrices $R$ and $S$ are derived through eigen decomposition on the covariance matrix $C$, which requires $O(d^{3})$ time and $O(d^{2})$ space using the Jacobi method. In Step 4, the time complexity of the MORBO algorithm in each iteration is $O(n_{QBS}^{3})$, where $n_{QBS}$ is the number of records in the QBS table. Since optimization of matrix $T$ requires $k$ iterations, the overall time complexity for Step 4 is bounded by $O(k\times n_{QBS}^{3})$. The space complexity for Step 4 is bounded by $O(d\times n_{QBS})$. After obtaining the optimal matrix $T$, the matrix multiplication $D\times T$ is performed, requires $O(N\times d^{2})$ time for computation and $O(N\times d)$ space for storing the results. Consequently, the overall time complexity of hyperspace transformation is bounded by $O(N)$ + $O(n_{QBS}^{3})$, and overall space complexity is bounded by $O(N)$ + $O(n_{QBS})$. As shown in Fig 10, we conduct experiments to evaluate the computational overhead of hyperspace transformation. The $T$ construction process includes steps 1 to 4, while the data transformation refers to the process of $D\times T$. The experimental results in Fig 10(a) demonstrate that we can further accelerate the hyperspace transformation by increasing parallelism. Fig 10(b) indicates that, as the dataset size increases, the hyperspace transformation process does not incur high computational overhead in Hudi’s parallel computing environment.
Table 5: Feature Selection and Feature Scaling Methods
Methods Type Invertibility Optimization based on Query Aware Clustering Improvement "Last-mile" Model Improvement Time Complexity Space Complexity
XGBoost(Chen and Guestrin, 2016) Feature Selection $O(KhNlogN)$ $O(KN)$
LDA(Xanthopoulos et al., 2013) Feature Selection $O(N)$ $O(N)$
PCA(Abdi and Williams, 2010) Feature Selection $O(N)$ $O(N)$
CCNF(Li et al., 2022a) Feature Scaling $O(KhNlogN)$ $O(KN)$
DTization(Islam, 2024) Feature Scaling $O(NlogN)$ $O(N)$
Ours Hyperspace Transformation Init_T:$O(N)$ Opt_T:$O(N)+O({n^{3}_{QBS}})$ Init_T:$O(N)$ Opt_T:$O(N)+O({n_{QBS}})$
  • a
    K is the number of tree.
  • b
    h is the depth of tree.
  • c
    Init_T(Initialized_T) represents the initialization process of matrix T in hyperspace transformation, which corresponds to Steps 1 - 3.
  • d
    Opt_T(Optimized_T) represents the initialization and optimization process of matrix T in hyperspace transformation which corresponds to Steps 1 - 4.
Next, we evaluate the effectiveness of hyperspace transformation. Our analytical framework includes two parts: theoretical analysis and experimental analysis. (1) Theoretical analysis: As shown in Table 5, we compare hyperspace transformation with several feature selection and feature scaling approaches, which share similarities with our hyperspace transformation process but also have notable differences. Feature selection is a dimensionality reduction technique aimed at selecting the most relevant features to enhance the performance and accuracy of machine learning algorithms (Dhal and Azad, 2022). Typical feature selection methods, such as XGBoost, PCA, and LDA, reduce dimensions and may eliminate crucial numerical features necessary for responding to specific queries, thus preventing the retrieval of certain information. On the other hand, feature scaling involves rescaling all features to a new scale, improving outcomes and speeding up computations in data processing and machine learning tasks (Alshaher, 2021). Typical feature scaling methods, such as CCNF and DTizatoin, can support invertibility by using an $O(N)$ space to store feature transformation information and could potentially replace our hyperspace transformation process. However, these feature scaling methods are not well-integrated with query-aware processes, cannot support clustering improvement, or facilitate high-quality "last-mile" model training. Additionally, they still exhibit higher time complexity compared to our approach. (2) Experimental analysis: We conduct experiments on these feature scaling methods compared with hyperspace transformation. As shown in Fig 11. The Optimized_T achieves the best average query time and recall rate, followed by Initialized_T, with the feature scaling methods CCNF and DTization performing worse than the hyperspace transformation, showing that our hyperspace transformation can better improve the efficiency and accuracy of multimodal data retrieval than other methods.
Figure 11: The average query times(a) and recall-time curves(b) of different methods in MQRLD.
@@ 5.2.3 Hyperspace Movement
Hyperspace movement further optimizes the feature vectors on the result of previous hyperspace transformation by moving similar data points in hyperspace together to get $\hat{D}$, where similar data can be placed in adjacent places or the same cluster. Thus it could benefit index construction and multimodal data retrieval.
To achieve this objective, we propose a new data movement method called Local Parallelized Gravitational Field (LPGF) which improves on an existing data movement method HIBOG (Li et al., 2021). HIBOG posits that in hyperspace, a data point is attracted by its nearest $K$ similar data points, causing it to move towards the resultant force direction of these attractions. Though HIBOG shows effectiveness in gathering similar data, it suffers the following drawbacks: (1) Computationally expensive: Finding the nearest $K$ data points for force calculation requires sorting all data points, which requires huge computation, especially for large datasets. Additionally, the parameter $K$ needs to be set and adjusted, further increasing the computation time. (2) Anomalies of movement: The calculation of forces between data points may not be suitable for all scenarios. Sometimes, this force can cause anomalies in the movement of tightly clustered data. (3) Weak parallel capability: The spatial range of the nearest $K$ neighbors is not predetermined, leading to the inability to perform grid partitioning and parallel processing of data. To overcome these shortcomings, we introduce the pipeline of LPGF, which consists of three steps to improve the above shortcomings, as shown in Fig 12.
Figure 12: Overview of hyperspace movement. Force computation of each point is calculated within radius R. All data points are diveded into n subspace, corresponding to n segments of Dataframe, and computed parallel in different nodes of cluster in Apach Hudi.
@@
Figure 13: Force calculation for each point is dependent on all points within its radius R, and the applying force varies for points at different distances.
@@ 5.2.4 Feature Representation Evaluation
The effectiveness of the feature representation is demonstrated through clustering performance, and clustering improves data layout’s separability and compactness. Meanwhile, superior data layout optimization can further enhance indexing performance, thereby benefiting efficient and effective queries. To validate this, we conduct three experiments. The first experiment compares the clustering performance before and after our feature representation process (Evaluation 1). The second and third experiments assess the query performance of learned indexes (Evaluation 2) and vector similarity indexes (Evaluation 3) on the optimized data layout, respectively.
Evaluation 1:  Clustering Performance Evaluation
In this evaluation, we show that the data layout after feature representation significantly enhances clustering performance. We compare five feature representation methods for enhancing clustering performance, which are T, HIBOG, LPGF, T + HIBOG, and T + LPGF. T represents Hyperspace Transformation. HIBOG and LPGF denote Hyperspace Movement. T + HIBOG as well as T + LPGF represent the combination of Hyperspace Transformation and Hyperspace Movement. The experiments are conducted using the real dataset Flame (Fu and Medico, 2007) and employ three well-known clustering methods, which are K-means, Agglomerative (Sokal and Michener, 1958), and Density Peaks Clustering (DPC) (Rodriguez and Laio, 2014), to represent diverse clustering paradigms. The selection of these methods is not meant to imply exclusivity but rather to demonstrate the generalization ability across various clustering algorithms. To comprehensively evaluate the impact of different representation methods on clustering performance, we use the Silhouette Coefficient, Calinski-Harabasz Index, and Normalized Mutual Information as experimental metrics. These metrics assess clustering effectiveness from different perspectives.
@@
  • b
    T: Apply hyperspace Transformation for feature enhancement.
  • c
    HIBOG: Apply HIBOG in hyperspace movement for feature enhancement.
  • d
    LPGF: Apply LPGF in hyperspace movement for feature enhancement.
@@
Table 6 presents the experimental results of different feature representation methods. Comparing the results of LPGF and HIBOG, we observe that LPGF significantly enhances clustering performance across various algorithms, suggesting that optimizations in gravitational field design in LPGF improve intra-class compactness and inter-class separation more effectively than HIBOG. Additionally, T + HIBOG and T + LPGF demonstrate that the hyperspace transformation matrix T not only enhances the discriminative features of multimodal data but also substantially improves clustering effectiveness.
@@
Evaluation 2:  Multi-dimensional Learned Index Query Performance Enhancement Evaluation
The core concept of the learned index is to use a learned model to narrow down queries to a "leaf node", followed by "last-mile" search within each leaf node for the final retrieval step. However, existing learned indexes face limitations in terms of both model accuracy and complexity. By optimizing the data layout, we improve clustering performance, which not only accelerates the "last mile" search but also improves both precision and recall. Additionally, the optimized data layout results in a better distribution of the data points, reducing the cost of model training, and thereby accelerating index construction. We design two experiments to validate the effectiveness of our proposed feature representation method in enhancing the performance of high-dimensional learned indexes. The first experiment shows that our approach enables the training of a "last mile" model which facilitates accelerated query processing and index construction, while the second evaluates its impact on shortening the query time and index construction time of current state-of-the-art multi-dimensional learned indexes. The experiments utilize a generated dataset comprising 2M records, with data distributed according to a Gaussian mixture distribution.
@@
In the first experiment, we predict the performance of the "last mile" model by computing the cumulative distribution function (CDF) of training labels for the "last mile" model, which we denote as keys. As shown in Fig 14, the key of a data point $P_{1_{n}}$ is computed as the sum of two distances. The first distance is from $P_{1_{n}}$ to its cluster’s centroid $C_{1}$, and the second distance is from $C_{1}$ to $C_{0}$ which is the barycenter of all clusters’ centroids. A smooth CDF curve suggests that the "last mile" model can effectively learn with high accuracy without the need for complex strategies, thereby reducing the query time as well as index construction time. We compare the CDF curves of four datasets, which are Original Dataset, HIBOG Optimized Dataset, LPGF Optimized Dataset, and T + LPGF Optimized Dataset. As shown in Fig 14, by comparing the smoothness of the CDF curves for each dataset, we find that T + LPGF achieves the highest improvement in the smoothness among all optimized methods. Compared to HIBOG, LPGF significantly smoothens the global CDF curve, because the force calculation method in LPGF reduces the probability of anomalous data movement, resulting in tighter data clustering. T + LPGF not only retains the advantages of LPGF but also highlights critical clustering features through matrix T, further enhancing the clustering capabilities. Overall, the T + LPGF method results in tighter multimodal data clustering, improving the smoothness of the CDF curve, and effectively reducing the complexity of the training process involved in index construction.
Figure 14: The CDF curves of different feature representation methods. Keys denote training labels of the "last-mile" search model. The CDF curve after T+LPGF is the smoothest, indicating that the "last-mile" model can be trained in a simpler, faster, and more accurate manner.
@@
In the second experiment, we demonstrate the effectiveness of our method in improving the performance of typical multi-dimensional learned indexes by comparing average query times and index construction times. We select five typical multi-dimensional learned indexes, which are ML, ZM, LISA, Flood, and LIMS. These learned indexes all share the characteristic that their query performance is significantly influenced by the smoothness of the CDF curve which directly impacts "last-mile" searching. As shown in Fig 15(a) and Fig 15(b), we compare the average query times and index construction times of different indexes on the original, the LPGF optimized, and the T + LPGF optimized datasets. The experimental results indicate that both LPGF and T + LPGF methods reduce query time and index construction time for these indexing methods, with T + LPGF exhibiting the most substantial improvements. On average, the query time of T + LPGF is 39.6% faster than that of the original data set, and the model training time is reduced by 56.7%.
Figure 15: The average query times(a) and index construction times(b) of different multi-dimensional indexes on original dataset and optimized dataset.
The average query times of ML, ZM, and LIMS have significantly improved because these index structures heavily depend on data distribution and clustering results, which confirms the effectiveness of our method in enhancing data clustering performance. LISA and Flood, being variants of grid indexes, benefit from our method’s ability to make data more tightly in each grid, resulting in faster indexing. For all the indexes, our method improves the data layout, leading to a smoother CDF curve that accelerates training time and ultimately reduces index construction time.
Figure 16: The average query times(a) and recall-time curves(b) of different vector similarity indexes on the original dataset and optimized dataset. Fig (a) shows that the query time of each index exhibits the most significant decrease after applying T+LGPF. Fig (b) shows that T+LGPF achieves the fastest attainment of 100% recall for each index.
Evaluation 3:  Vector Similarity Index Query Performance Enhancement Evaluation
Vector similarity index identifies similar data points by comparing their distances in space. Our feature representation process brings similar data points closer together in the space, leading to better clustering results and improved performance of vector similarity indexes. To demonstrate the effectiveness of our method in enhancing vector index performance, we compare the average query time and the Time-Recall curve. We select four vector indexes: DB-LSH, IVF, HNSW, and LSH. The performance of these indexes is significantly influenced by the data distribution, with datasets exhibiting clear clustering features showing substantial improvements in index performance. As shown in Fig 16(a) and Fig 16(b), we compare the average query time and recall rate for queries on the original, LPGF-optimized, and T + LPGF-optimized datasets. From the comparison results, we observe that all vector similarity indexes show the greatest improvements on the T + LPGF-optimized dataset, both in terms of average query time and the time required to achieve 100% recall. T + LPGF reduces the query time by an average of 42.8% compared to the original dataset, and the time to reach 100% recall is shortened by 55.8%. IVF shows the most significant improvement, as the T + LPGF method groups similar data points within the same cluster as well as clarifies the boundaries of different clusters. This enables the recall rate to reach 100% in a shorter time. DB-LSH, a hybrid index, relies on the hash function to partition data based on the hash values. The T + LPGF method improves this by tightly clustering similar data points, effectively reducing the error bucket rate caused by hash collisions. Additionally, the advantage of T + LPGF makes the centroids of different clusters more distinct. Using these centroids to build the HNSW graph structure significantly reduces the complexity of each layer, enhancing the data navigation efficiency.
@@ 6 High-dimensional Learned Index
To fully leverage the advantages of the "Multimodal Data Feature Representation" introduced in Section 5 and further enhance multimodal data retrieval efficiency, we propose a high-dimensional learned index. Unlike multi-dimensional learned indexes like Flood and Tsunami, etc. which are restricted to multiple numeric queries, or ML and LIMS, etc. which only support single numeric queries and vector queries, our index is designed for rich hybrid queries and performs better on both numeric and vector queries than these indexes. Compared to vector similarity indexes such as FLANN and HNSW, etc. which focus on vector queries on high-dimensional data, our index provides more versatile query options and offers better efficiency and effectiveness on large datasets. Our index hierarchically organizes clusters and their sub-clusters to form a cluster tree, where the index of the leaf nodes is established through a "last-mile" training approach. Furthermore, the index’s inner structure can be optimized through our query-aware mechanism. An overview of our index is illustrated in Fig 17, detailing both the construction and optimization processes. During the construction phase, divisive hierarchical clustering is employed to form clusters at various levels, corresponding to different levels within the tree structure, with the leaf nodes representing the most granular, indivisible sub-clusters, thereby substantially enhancing the precision of the "last-mile" training for the index. Concurrently, the optimization phase employs a query-aware mechanism to refine the internal structure of the index, improving its adaptability to diverse query demands. The detailed processes of index construction and optimization are elaborated in Section 6.1 and Section 6.2, respectively.
Figure 17: The index construction process consists of divisive hierarchical clustering, cluster tree building, and query-aware optimization.